1/1

Lord George Gordon Byron ,

Sun of the Sleepless

Sun of the Sleepless! melancholy star! Whose tearful beam glows tremulously far, That show’st the darkness thou canst not dispel, How like art thou to Joy remembered well! So gleams the past, the light of other days, Which shines but warms not with its powerless rays: A night-beam Sorrow watcheth to behold, Distinct, but distant – clear – but, oh, how cold!

Verkinto de tiu ĉi Angla poemo estas Lord George Gordon Byron (∗1788-01-22 – †1824-04-18).

Arg-1101-2219 (2014-12-11 17:39:08)

Tiu ĉi poemo troveblas en http: // www. poetryloverspage. com/ poets/ byron/ sun_of_sleepless. html . Pri la verkinto vidu la retejon http: // eo. wikipedia. org/ wiki/ George_Byron .

http://www.poezio.net/version?poem-id=1101&version-id=,2219